# Surrogate modeling of Cellular-Potts Agent-Based Models as a segmentation task using the U-Net neural network architecture

### v3

@@ 4 Discussion
In summary, we demonstrate the applicability of convolutional neural networks classically applied to segmentation tasks to perform as surrogates of Cellular-Potts agent-based models inclusive of PDE-based diffusive fields. Our surrogate predicts the simulation configuration significantly faster than native Cellular-Potts algorithm execution and can replicate emergent behaviors present but not explicitly defined in the Cellular-Potts model. Overall, this work highlights the potential of deep learning approaches to accelerate the evaluation of Cellular-Potts models, offering a promising direction for improving computational efficiency for complex biological simulations.
@@ Acknowledgements
Code for this work will be made publicly available at https://github.com/tc2fh/CPM_UNet_Surrogate prior to publication.

### v4

@@ Conclusion
In summary, we demonstrate the applicability of convolutional neural networks classically applied to segmentation tasks to perform as surrogates of Cellular-Potts agent-based models inclusive of PDE-based diffusive fields. Our surrogate predicts the simulation configuration significantly faster than native Cellular-Potts algorithm execution and can replicate several multicellular emergent tissue-scale behaviors not explicitly encoded in the behaviors of individual CPM agents. While the model is effective at shorter timescales (up to 3 recurrent evaluations), its predictions diverge from the mechanistic CPM simulation over longer times. Overall, this work highlights the potential of deep learning approaches to accelerate the evaluation of Cellular-Potts models, offering a promising direction for improving computational efficiency for complex biological simulations.
Supporting information
S1 Movie
Surrogate model vascular network configuration prediction with recurrent evaluation. Ground truth (left) from the Cellular-Potts model is compared with the surrogate model prediction (right) given the same initial configuration.
S2 Movie
Surrogate model diffusive field prediction with recurrent evaluation. Ground truth (left) from the Cellular-Potts model is compared with the surrogate model prediction (right) given the same initial configuration.
@@ Acknowledgements
Code to generate data and specify models used in this work is publicly available at https://github.com/tc2fh/CPM_UNet_Surrogate. This code is also published on Zenodo at https://zenodo.org/records/15399533[38] with a Creative Commons Attribution 4.0 International license.